OpenStudy (anonymous):

There are 48 students in a school play. the ratio of boys to girls is 5:7. How many more girls than boys are in the play? Explain how you found your answer.

OpenStudy (anonymous):

well, for some integer k, there must be 5k+7k = 48 students

OpenStudy (anonymous):

since 7k - 5k is the difference .... id focus on finding k

OpenStudy (a1234):

I think would need to do 48/12, which is 4. 4*5 = 20 4*7 = 28

OpenStudy (anonymous):

Thanks ^

OpenStudy (a1234):

There are 20 boys and 28 girls.